OpenStudy (anonymous):

A spherical balloon is expanding at a constant rate of 32(pi) cm^3/s. The balloon is intially fully deflated. After 9 seconds, the volume of the balloon is 288(pi) cm^3. At this instant, find the rate of increase of the radius of the balloon. Answer given: 2/9 cm/s

OpenStudy (anonymous):

We know the rate of expansion of the balloon. \[\frac{ dV }{ dt}=32\pi \] We want to find \[\frac{ dr }{ dt }\] so we try to equate the two and see what's missing \[\frac{ dr }{ dt }=\frac{ dV }{ dt }\times? \] We easily see that the missing factor is \[\frac{ dr }{ dV }\] to find this, we can find the the rate of change of volume with respect to radius from the volume of a sphere \[V=\frac{ 4 }{ 3 }\pi r ^{3}\] \[\frac{ dV }{ dr }=4 \pi r ^{2}\] Flipping it we get \[\frac{ dr }{ dV }=\frac{ 1 }{ 4 \pi r ^{2} }\] Now plugging it i gives \[\frac{ dr }{ dt }=32 \pi \times \frac{ 1 }{ 4 \pi r ^{2} }=\frac{ 8 }{ r ^{2} }\]

OpenStudy (anonymous):

At that instance, we know the volume is 288 Pi. So we can find r from the formula for the volume of a sphere. r is found to be 6, Thus, dr/dt = 2/9

OpenStudy (anonymous):

@Rapidknight thank you so so much!

OpenStudy (anonymous):

You're welcome =)